If i travel 1000 miles at 62mph how long would it take?
Answer by jim_thompson5910(35256)

D = r*t


1000 = 62*t


1000/62 = t


16.12903 = t


t = 16.12903


It will take approximately 16.12903 hours
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------